solve
(x-6)(x+4)=11

Answer by mukhopadhyay(490)

(x-6)(x+4)=11
=>x^2-2x-24=11
=>x^2-2x-35=0
=>x^2-7x+5x-35=0
=>(x-7)(x+5)=0
=>x=7 or x=-5
Answer: x=7 or x=-5